OpenStudy (anonymous):

OK gonna need some help... Don't remember this! Find sin 540° + 4(tan150°)(cot120°)

OpenStudy (anonymous):

sin 540 = sin 180 = 0 tan 150 = -1 / sqrt(3) cot 120 = -1 / sqrt(3)

OpenStudy (anonymous):

\[0 + 4 \times ( {-1 \over \sqrt{3}} \times {-1 \over \sqrt{3}} )\] \[4 \times { { -1 \times -1 } \over \sqrt3 \times \sqrt3 }\] \[4 \times { 1 \over 3 } = { 4 \over 3}\]

OpenStudy (anonymous):

Is that how you do it? crazy I would have never got that in a million years.

OpenStudy (anonymous):

Simple, i calculate the results for the 3 different goniometric expressions, then feed them into the given formula. Remember, with math, it's not enough to remember something only for a test. You'll need everything you learn in subsequent tests.

OpenStudy (anonymous):

thought tan = 150 would be 30 degrees which is sqrt(3)/3 and cot = 120 would be 60 degrees which is sqrt(3)/3

OpenStudy (anonymous):

wouldn't they both be negative as well?

OpenStudy (anonymous):

yep, they are, but a negative times a negative is a positive

OpenStudy (anonymous):

since sqrt(3)/3 x sqrt(3)/3 = 3/9 and 0 + 4(3/9) = 7/9? So it would be 7/9

OpenStudy (mertsj):

\[\frac{4}{1}\times\frac{3}{9}=\frac{12}{9}=\frac{4}{3}\]